OpenStudy (3mar):

Solve (a) x^2+2x-8=0 and (b) 3x^2 - 11x -4=0 by using the quadratic formula.

OpenStudy (baayen):

Do you know what the quadratic formula is?

OpenStudy (3mar):

I know factorization!

OpenStudy (baayen):

Okay good, but this problem is asking you to use the quadratic formula. Does \[-b \pm \sqrt{b ^{2}-4ac} \over 2a \] look famillar?

OpenStudy (3mar):

Yes, then?

OpenStudy (baayen):

x^2+2x-8=0, so looking at this equation can you tell me what is a b and c? hint, standard form is ax^2+ bx+c=0

OpenStudy (3mar):

a=1 b=2 c=8

OpenStudy (baayen):

Okay good, now plug those into the formula I gave you earlier

OpenStudy (3mar):

Yes. and simplify and get two values. is not it?

OpenStudy (baayen):

No that's correct, do you know what those two values are?

OpenStudy (3mar):

did not get them yet

OpenStudy (baayen):

Mind showing me what you got after you plugged them into the quadratic formula?

OpenStudy (3mar):

gig you get them?

OpenStudy (baayen):

Are you asking me to just give you the answers? lol

OpenStudy (3mar):

Thanks a lot. I got the answer.

OpenStudy (baayen):

Was that sarcasm? Can't tell sorry

OpenStudy (3mar):

Excuse me??